calculate the number of atoms present in 100g of Ag
1
Expert's answer
2020-11-23T06:50:07-0500

Quantity of Ag:

100 g / (107.87 g/mol) = 0.927 mol

Number of atoms:

0.927 mol * (6,022*1023 mol-1) = 5.582*1023

Answer: 5.582*1023 atoms